OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. 6, 12, 18

OpenStudy (anonymous):



HanAkoSolo (jamierox4ev3r):

THats right @GreenBeanDemphsey :) And just so you know, the Pythagorean Theorem is \(A^2+B^2=C^2\)

HanAkoSolo (jamierox4ev3r):

can you figure this out now based on the info. I gave you @GreenBeanDemphsey ?

HanAkoSolo (jamierox4ev3r):

hellloooo you there? O_O

OpenStudy (anonymous):

sorry i was cooking dinner and thank you!

HanAkoSolo (jamierox4ev3r):

no problem! So can you solve for me ? :D

OpenStudy (anonymous):

i just need to make sure i was right but if you want me to i will

HanAkoSolo (jamierox4ev3r):

yeah i want to make sure you understand! :)

OpenStudy (anonymous):

The question is basically asking if this equation is true: \[6^2 + 12^2 = 18^2?\]Which simplifies into\[36+144=324\]or\[180=324\]Since 180 is not in fact equal to 324, it can be concluded that this equation is false, so 6,12, and 18 do not fit the Pythagorean theorem.

HanAkoSolo (jamierox4ev3r):

@slottedpig I was asking @GreenBeanDemphsey to solve it... -.- please do not provide answers, let the user find it themselves!!!

OpenStudy (anonymous):

what i thought it was asking is 18 was a possiable hypthenuse? 0.o

HanAkoSolo (jamierox4ev3r):

no the problem is actualy asking whether the measures could become a right triangle....